OpenStudy (anonymous):

Solve 8(t+2)-3(t-4)=6(t-7)+8 in steps?

OpenStudy (anonymous):

8(t+2)-3(t-4)=6(t-7)+8 <=> 8t + 16 - 3t + 12 = 6t - 42 + 8 <=> 5t + 28 = 6t - 34 <=> 6t - 5t = 28 + 34 <=> t = 62

OpenStudy (anonymous):

Thanks